If y=sin2ax+cosax,prove that yn=a^n{1+(-1) sin2ax}^1/2


1
Expert's answer
2021-10-01T15:52:01-0400

"\begin{aligned}
y &=\sin a x+\cos a x \\
y_{n} &=a^{n} \sin \left\{\frac{n \pi}{2}+a x\right\}+b^{n} \cos \left\{\frac{n \pi}{2}+a x\right\} \\
&=a^{n}\left(\sin \left\{\frac{n \pi}{2}+a x\right\}+\cos \left\{\frac{n \pi}{2}+a x\right\}\right)
\end{aligned}\\
 
Since (\sin a x+\cos a x)=(1+\sin 2 a x)^{\frac{1}{2}} , then\\
 
\begin{aligned}
y_{n} &=a^{n}\left\{1+\sin 2\left(\frac{n \pi}{2}+a x\right)\right\} \\
&=a^{n}\left\{1+\sin \left(\frac{2 n \pi}{2}+2 a x\right)\right\}^{\frac{1}{2}} \\
&=\left\{1+(-1)^{n} \sin 2 a x\right\}^{\frac{1}{2}}
\end{aligned}"